# An angular-momentum preserving dissipative model for the point-mass $N$-body problem

### 12 Mar 2026 v1

@@ Abstract
A simple mathematical model emulating energy dissipation due to tidal effects is proposed. In this model, forces acting between masses remove energy but preserve the total angular momentum of the system. We study the effect of such forces on a particular family of orbits, namely those corresponding to central configurations, and show that a specific dependence on the mutual distances between the bodies leads to homographic equations equivalent to those of the two-body problem with dissipation. We then describe in detail the topology of solutions of the dissipative two-body system via Poincaré compactification. Finally, we present equations averaged over Keplerian motion showing no influence of the dissipation on periapsis precession.
@@
Celestial Mechanics, $N$-body problem, Dissipative systems, Transient dynamics
@@ 1 Introduction
In this context, Smale [Smale] proved that for any gravitational system of point-masses, the critical points of the energy function (1) restricted to a constant angular momentum (4) hypersurface correspond to relative equilibrium configurations, i.e.,
@@
In the presence of tides, which dissipates energy but not angular momentum, many dynamical conclusions can be derived from the results of Smale and Moeckel. For instance, for $N\geq 2$, energy dissipation leads to two possible physical scenarios: either bodies collide and/or move apart and ultimately become isolated masses; or, after successive collisions and escapes, pairs of masses remain in circular motion, where no further energy can be dissipated. See [Scheeres_2002, Scheeres_2012] for details. The combination of Smale’s and Moeckel’s results assert that, despite the existence of relative equilibria, their instability prevents systems with more than two bodies to remain in such states indefinitely.
@@
Even though these models correspond to relevant and realistic scenarios, they remove both energy and angular momentum from the system. By contrast, at planetary scale, the most prominent dissipative effect is gravitational in origin, arising from tides [Goldreich, Meyer], which preserves angular momentum.
@@
We remark that a force similar, though not identical, to the one we propose has already appeared in a PhD thesis by Chong Ai under the supervision of Mark Levi [Chong_Ai]. Since their force dissipates energy while preserving angular momentum, their qualitative results for the dissipative Kepler problem are similar to ours. They also studied dissipative effects in the restricted three-body problem, an investigation that we are also pursuing. To our knowledge, the exact force expression we propose, which under dissipation leaves invariant a surface that in the conservative case is generated by the homographic solutions, has not been proposed before.
@@ 2 Model
where $\hat{\textbf{r}}_{ij}=\frac{\textbf{r}_{i}-\textbf{r}_{j}}{|\textbf{r}_{i}-\textbf{r}_{j}|}\in\mathbb{R}^{3}$ is the versor connecting the $i$-th and $j$-th bodies. The dissipation strength $\alpha$ is constant and has units of $\left[L^{d-3}T\right]$. Physically, it reflects rheological and viscoelastic effects, carrying the dependence on the body’s Love number (deviation from a spherical shape), equatorial radius, and tidal torque delay  [^1] In principle, the dissipation constant $\alpha$ may vary from body to body, so that $\textbf{f}^{\,ij}_{\textrm{diss}}(\dot{\textbf{r}}_{ij},\textbf{r}_{ij})=-\alpha_{i}\alpha_{j}\frac{Gm_{i}m_{j}}{|\textbf{r}_{ij}|^{d}}\left(\dot{\textbf{r}}_{ij}\cdot\hat{\textbf{r}}_{ij}\right)\hat{\textbf{r}}_{ij},$ In a forthcoming paper, focusing on the three-body problem, we use this formulation. Here we assume a uniform $\alpha$ for all bodies. This distinction is irrelevant in the case of two bodies. . Since the current study prioritize a mathematical approach, we treat the radial dependence exponent $d$ as a free parameter, as it will be discussed further in the text. Physically, a direct comparison to Mignard’s model would yield $d=8$.
@@
where the positions $\textbf{r}_{i}\in\mathbb{R}^{3}$ for each mass $m_{i}$ are relative to the center of mass.
@@ 2.1 Energy dissipation
In fact, the only case where dissipation can cease indefinitely is for two bodies. For $N=2$, energy reaches its minimum at the circular relative equilibrium and remains there. Whenever other bodies are present, even if the movement is locally circular, the presence of any remaining body keeps $\mathcal{D}\neq 0$ at all times, unless they all rotate as a rigid body.
@@ 3 Central configurations
Given a set of $N$-bodies with masses $m_{1},\dots,m_{N}$, a CC is an arrangement where the force on each body constantly points towards the center of mass of the system, that is
@@
stores the masses for each body, with $\textbf{M}_{i}=m_{i}\mathbf{I}_{3}$ (or $m_{i}\mathbf{I}_{2}$), for i=1,,Ni=1,...,N. The vector $\textbf{c}\in\mathbb{R}^{3N}$ is a stacked vector of the center of mass $\textbf{c}_{0}$, such that
@@ 3.1 Planar case
where $s(t)\in\mathbb{R}$ is the scaling factor, $\textbf{Q}(\theta(t))$ is a rotation matrix perpendicular to the plane, such that each body is rotated equally by $\theta(t)$
@@
Over a homographic configuration ($\textbf{r}_{h},\dot{\textbf{r}}_{h}$) satisfying (17) the Rayleigh dissipation function (10) can be evaluated
@@
$\begin{split}&\frac{d\left(\textbf{r}-\textbf{c}\right)}{dt}=\left(\dot{s}(t)\textbf{Q}(\theta(t))+s(t)\textbf{Q}^{\prime}\dot{\theta}(t)\right)\left(\textbf{r}_{0}-\textbf{c}_{0}\right)\\ &\frac{d^{2}\left(\textbf{r}-\textbf{c}\right)}{dt^{2}}=\left(\ddot{s}(t)\textbf{Q}(\theta(t))+2\dot{s}(t)\textbf{Q}^{\prime}\dot{\theta}(t)+s(t)\textbf{Q}^{\prime\prime}\dot{\theta}^{2}(t)+s(t)\textbf{Q}^{\prime}\ddot{\theta}(t)\right)\left(\textbf{r}_{0}-\textbf{c}_{0}\right)\end{split}$
@@
It was also used that rotation matrices preserve inner products and that, being restricted to the plane they commute, implying
@@ 3.2 The special case $d=3$
$\begin{split}&\nabla_{\dot{\textbf{r}}_{i}}\mathcal{D}(\textbf{r}_{h},\dot{\textbf{r}}_{h})=\alpha\dot{s}s^{-3}\textbf{q}\nabla_{\textbf{r}_{i}}U(\textbf{r}_{0}).\end{split}$ (21)
From a mathematical perspective, it is interesting to consider the case $d=3$ as it produces an additional symmetry between the dissipative and conservative terms. This equivalence allows us to obtain the equation for the homographic parameters ($s(t),\theta(t)$) in the presence of dissipation.
Substituting the homographic form (17) into the stacked equations of motion (12) for all bodies and multiplying by $\textbf{Q}^{-1}$ gives
@@
$\nabla_{\textbf{r}_{0}}U(\textbf{r}_{0})+\lambda\textbf{M}\left(\textbf{r}_{0}-\textbf{c}_{0}\right)=0.$
Simultaneously, since J is a right-angle rotation, it renders $\textbf{J}(\textbf{r}_{0}-\textbf{c}_{0})$ orthogonal to $(\textbf{r}_{0}-\textbf{c}_{0})$. This splits equation (22) into terms parallel and orthogonal to $(\textbf{r}_{0}-\textbf{c}_{0})$, yielding a system of equations for the coefficients
$\begin{cases}\ddot{s}-s\dot{\theta}^{2}=-\lambda\left(s^{-2}+\alpha\dot{s}s^{-3}\right)\\ s\ddot{\theta}+2\dot{s}\dot{\theta}=0\end{cases}$ (23)
Equation (23) is a dissipative Kepler-like system in the homographic parameters, with $s(t)$ playing the role of the radial coordinate and $\theta(t)$ the angular position. Due to this equivalence, the next section treats the two-body problem under dissipation with $d=3$, providing solutions applicable both to the two-body case and to generic CCs for arbitrary $N$.
@@ 4 Two bodies
We describe the two-body system in polar coordinates $(r,\theta)$, where $v=\dot{r}$ denotes the radial velocity. The equations of motion take the first-order form
@@
$\mathcal{H}=\frac{v^{2}}{2}+\frac{C^{2}}{2}-\frac{\gamma}{r}\qquad\text{and}\qquad\mathcal{D}=\frac{\alpha}{2}\frac{\gamma}{r^{3}}v^{2},$ (25)
and the (conserved) angular momentum $C$ defined by
@@
To solve the movement in equations (24), we proceed to a Poincaré compactification of the phase space in a series of transformations, following Dumortier et al. [Dumortier_book] (Chapter 5.1). The idea relies on mapping the infinite phase space $(r,v)\in[0,\infty)\times\mathbb{R}$ into a sphere, and analyze the dynamics at infinity via stereographic charts (planar maps) covering the sphere. This allows us to understand the asymptotic behavior of solutions as they approach infinity in the original coordinates.
@@ 4.1 Poincaré compactification
$\begin{split}a^{c}_{\pm}=\frac{1}{2}\left(-\alpha\gamma\pm\sqrt{\alpha^{2}\gamma^{2}-\frac{4C^{6}}{\gamma^{2}}}\right),\end{split}$
@@
Hence, the equilibrium $(r_{e},v_{e})$ has an eigenvalue $a^{e}=-\gamma\alpha$ and a stable manifold coinciding with the axis $r=0$. The other eigenvalue is zero and the associated center-manifold is tangent to the vector $(\alpha,-1)$ (Fig. 1 (C)).
The physical meaning of the equilibrium $(r_{e},v_{e})$ is a direction towards which all orbits with initials conditions $(r(0),v(0))$, for fixed $v(0)$ and $r(0)>0$ arbitrarily small, must approach before entering the circularization spiral towards $(r_{c},v_{c})$, as shown in Fig. 1 (B). In this figure, streamlines density and velocity close to $r=0$ are much smaller than in the original dissipative Kepler problem (eq. 24).
Figure 1: Phase space portraits for the regularized two body system (eq. 27) for $C=\gamma=1$. (A) The conservative case ($\alpha=0$) showing elliptic (blue), parabolic (green), and hyperbolic (orange) orbits. (B) The system with dissipation ($\alpha=0.5$) showing a captured orbit (blue), the weak-unstable-center manifold (red) from the equilibrium $(r_{e},v_{e})=(0,2)$, and an escaping orbit (orange); the green dot marks the circular equilibrium $(r_{c},v_{c})=(1,0)$. (C) Zoom of (B) near $(r_{e},v_{e})$; stable branches along $r=0$ are shown in blue.
@@
$(r,v)\longmapsto(y,x)\quad\text{where}\quad r=\frac{1}{y},\;\;v=\frac{x}{y}$
@@
In the new coordinates, equation (27), after a new multiplication by $y^{3}$ becomes
@@
$(y_{\infty},x_{\infty})=(0,0)$
corresponding to the limit as $r\to\infty$ with $v=$ constant. As the linear approximation of (30) at $(0,0)$ is null, $(y_{\infty},x_{\infty})$ is a degenerate equilibrium. Thus, all unbounded solutions of the Kepler problem ($\alpha=0$) are homoclinic to $(y_{\infty},x_{\infty})=(0,0)$ (see Fig. 2 (A)).
@@
Locally, the degenerate equilibrium $(y_{\infty},x_{\infty})$ has a characteristic curve $\Gamma_{s}$ approaching it for positive time, and a curve $\Gamma_{u}$ approaching it for negative time. We can also divide the phase space xxyy into three sectors: a hyperbolic sector $S_{h}$, an attracting sector $S_{a}$, and a repelling sector $S_{r}$ (see Fig. 2 (C)).
To provide a local approximation to $\Gamma_{u(s)}$, we blow up the singularity (following [Arnold_book], Sec. 1.2), by partially reverting the transformation $(y,x)\mapsto(y,v)$, where $v=\frac{x}{y}$. System 30 then becomes (after a division by $y^{2}$):
[…]
which is no longer singular at the origin, as $\left.(\dot{y},\dot{v})\right|_{(0,0)}=(0,-\gamma)$. Thus, for $|v|$ small
$\frac{dy}{dv}=\frac{v}{\gamma(1+\alpha yv)-C^{2}y}$
which has Taylor expansion near ($0,0$) as
[…]
@@
$\Gamma_{u(s)}=x_{+(-)}=\pm\sqrt{2\gamma}\,y^{\frac{3}{2}},$ (32)
@@
For $\alpha=0$, the vector field (30) has the first integral $\mathcal{H}=\frac{y^{3}\left(C^{2}y-2\gamma\right)+x^{2}}{2y^{2}}$ (from eq. 25), and both characteristic curves $\Gamma_{u}$ and $\Gamma_{s}$ are determined by the equation $\mathcal{H}=0$. One can check that this leads to the solution in equation (32).
In the Kepler problem, orbits with positive energy are asymptotically free, with $\mathcal{H}\to\frac{v^{2}}{2}$ as $t\to\pm\infty$. Scattering orbits in the dissipative or conservative scenarios are those for which the following two limits exist
@@
In order to compute the asymptotic shape of a scattering orbit close to a point at infinity, one must solve
$\mathcal{H}=\frac{y^{3}\left(C^{2}y-2\gamma\right)+x^{2}}{2y^{2}}=\frac{v^{2}}{2}\quad\text{or}\quad y^{3}\left(C^{2}y-2\gamma\right)+x^{2}=v^{2}y^{2}$
@@
$x=\pm vy+\mathcal{O}(y^{3/2})$
Therefore, a scattering orbit leaves the point at infinity tangent to the line $x=-v_{\psi}y$, and arrives at the same point tangent to the line $x=v_{\omega}y$. In the presence of dissipation, $|v_{\omega}|<|v_{\psi}|$.
Besides, the equation for the characteristic curves $\Gamma_{u(s)}$, $x=\pm\sqrt{2\gamma}y^{\frac{3}{2}}$ implies that scattering orbits arrive and leave $(x_{\infty},y_{\infty})$ tangent to the yy–axis. Therefore, as expected, scaterring orbits must both leave the point at infinity by the repelling sector $S_{r}$ and arrive at the point at infinity in the attracting sector $S_{a}$ (see Fig. 2 (C)).
@@
$(r,v)\mapsto(w,z)\quad\text{where}\quad w=\frac{r}{v},\quad z=\frac{1}{v}$
In these coordinates, $z=0$ corresponds to $v=\infty$.
After a convenient time reparametrization (again multiplying by $z^{3}$) the vector field (27) in the new chart becomes
@@
$(w,z)=(w_{v\infty},z_{v\infty})=(0,0)$
representing the limit as $v\to\infty$ with r=r=constant. As the linear approximation to the vector field (33) at $(0,0)$ is null, $(w_{v\infty},z_{v\infty})$ is a degenerate equilibrium. The circular equilibrium $(r_{c},v_{c})$ is mapped to infinity in this chart.
In Fig. 3 we show phase portraits for the system with and without dissipation. The maping of orbits from $(r,v)$ to $(w,z)$ is not fully clear. Bounded orbits in $(r,v)$ are mapped to a pair of curves in $(w,z)$ that extend to infinity as any periodic orbit requires $v=0$ that maps to $z,w\to\infty$. Consequently, a single initial condition ($w_{0},z_{0}$) covers only one branch and is not able to reproduce the whole mapping of a bounded ($r,v$) orbit. The full periodic curve is recovered by gluing the two branches from the opposite quadrants ($w>0,z>0$) and ($w<0,z<0$). The remaining quadrants, ($w>0,z<0$) and ($w<0,z>0$) do not correspond to physical orbits, as in these regions $r<0$ (see Fig. 3 (A)).
For $w=0$ we have $\dot{w}=0$ and $\dot{z}=\dot{z}(z)$, so the $z$-axis is invariant. In the presence of dissipation ($\alpha>0$), the equilibrium $(w_{e},z_{e})$ lies over the $z$-axis and there is a heteroclinic connection with $(w_{v\infty},z_{v\infty})$. The unstable manifold of $(w_{e},z_{e})$ (pink branches in Fig. 3 (B)) separates orbits of close approach with limited velocity (above the unstable branch), from those that can be arbitrarily close and reach infinite velocities (below the unstable branch).

### 06 Apr 2026 v2

@@ Abstract
A simple mathematical model emulating energy dissipation due to tidal effects is proposed. In this model, forces acting between masses remove energy but preserve the total angular momentum of the system. We study the effect of such forces on the particular family of orbits in central configurations, and show that a specific dependence on the mutual distances between the bodies leads to homographic equations equivalent to those of the two-body problem with dissipation. We then describe in detail the topology of solutions of the dissipative two-body system via Poincaré compactification. Finally, we present equations averaged over Keplerian motion showing no influence of the dissipation on periapsis precession.
@@
Celestial Mechanics, $N$-body problem, Dissipative systems
@@ 1 Introduction
In this context, Smale [Smale] proved that for any gravitational system of point-masses, the critical points of the energy function (1) restricted to a constant angular momentum hypersurface (4) correspond to relative equilibrium configurations, i.e.,
@@
In the presence of tides, which dissipate energy but not angular momentum, a distinct conclusion can be derived from the results of Smale and Moeckel. For $N\geq 2$, energy dissipation leads to two possible physical scenarios: either bodies collide and/or move apart and ultimately become isolated masses; or, after successive collisions and escapes, pairs of masses remain in circular motion, where no further energy can be dissipated (see [Scheeres_2002, Scheeres_2012] for details). The combination of Smale’s and Moeckel’s results assert that, despite the existence of relative equilibria, their instability prevents systems with more than two bodies to remain in such states indefinitely.
@@
Even though these models correspond to relevant and realistic scenarios, they remove both energy and angular momentum from the system. By contrast, at planetary scale, the most prominent dissipative effect is gravitational in origin, arising from tides, which preserves angular momentum [Goldreich, Meyer].
@@
We remark that a force similar, though not identical, to the one we propose has already appeared in a PhD thesis by Chong Ai under the supervision of Mark Levi [Chong_Ai]. Since their force has the same physical features as the one we propose here, their qualitative results for the dissipative Kepler problem are similar to ours. They also studied dissipative effects in the restricted three-body problem, an investigation that we are also pursuing. To our knowledge, the exact force expression we propose, which under dissipation leaves invariant a manifold that in the conservative case is generated by the homographic solutions, has not been proposed before.
@@ 2 Model
where $\hat{\textbf{r}}_{ij}=\frac{\textbf{r}_{i}-\textbf{r}_{j}}{|\textbf{r}_{i}-\textbf{r}_{j}|}\in\mathbb{R}^{3}$ is the versor connecting the $i$-th and $j$-th bodies. The dissipation strength $\alpha$ is constant and has units of $\left[L^{d-3}T\right]$. Physically, it reflects rheological and viscoelastic effects [^1] In principle, the dissipation constant $\alpha$ may vary from body to body, so that $\textbf{f}^{\,ij}_{\textrm{diss}}(\dot{\textbf{r}}_{ij},\textbf{r}_{ij})=-\alpha_{i}\alpha_{j}\frac{Gm_{i}m_{j}}{|\textbf{r}_{ij}|^{d}}\left(\dot{\textbf{r}}_{ij}\cdot\hat{\textbf{r}}_{ij}\right)\hat{\textbf{r}}_{ij},$ In a forthcoming paper, focusing on the three-body problem, we use this formulation. Here we assume a uniform $\alpha$ for all bodies. This distinction is irrelevant in the case of two bodies. . Since the current study focuses on a mathematical approach, we treat the radial dependence exponent $d$ as a free parameter, as it will be discussed further in the text. A direct comparison to Mignard’s model would yield $d=8$.
@@
for $\textbf{r}_{i}\in\mathbb{R}^{3}$ as the position of each mass $m_{i}$ relative to the center of mass.
@@ 3 Central configurations
Given a set of $N$-bodies with masses $m_{1},\dots,m_{N}$, a CC is an arrangement where the force on each body points towards the center of mass of the system, that is
@@
stores the masses for each body, with $\textbf{M}_{i}=m_{i}\mathbf{I}_{3}$ (or $m_{i}\mathbf{I}_{2}$), for i=1,,Ni=1,...,N, and where $\mathbf{I}_{3}$ ($\mathbf{I}_{2}$) is the identity matrix in $\mathbb{R}^{3}$ ($\mathbb{R}^{2}$). The vector $\textbf{c}\in\mathbb{R}^{3N}$ is a stacked vector of the center of mass $\textbf{c}_{0}$, such that
@@ 3.1 Planar case
where $s(t)\in\mathbb{R}$ is the scaling factor, and $\textbf{Q}(\theta(t))$ is a rotation matrix about the axis perpendicular to the plane, rotating each body uniformly by $\theta(t)$
@@
Over a homographic configuration ($\textbf{r}_{h},\dot{\textbf{r}}_{h}$) satisfying (17) the Rayleigh dissipation function (10) can be evaluated as
@@
$\begin{split}&\frac{d\left(\textbf{r}-\textbf{c}\right)}{dt}=\left(\dot{s}(t)\textbf{Q}(\theta(t))+s(t)\textbf{Q}^{\prime}\dot{\theta}(t)\right)\left(\textbf{r}_{0}-\textbf{c}_{0}\right)\\ &\frac{d^{2}\left(\textbf{r}-\textbf{c}\right)}{dt^{2}}=\left(\ddot{s}(t)\textbf{Q}(\theta(t))+2\dot{s}(t)\textbf{Q}^{\prime}\dot{\theta}(t)+s(t)\textbf{Q}^{\prime\prime}\dot{\theta}^{2}(t)+s(t)\textbf{Q}^{\prime}\ddot{\theta}(t)\right)\left(\textbf{r}_{0}-\textbf{c}_{0}\right)\end{split}$ (21)
@@
It was also used that rotations preserve inner products and that, when restricted to the plane, they commute, implying
@@ 3.2 The special case $d=3$
$\begin{split}&\nabla_{\dot{\textbf{r}}_{i}}\mathcal{D}(\textbf{r}_{h},\dot{\textbf{r}}_{h})=\alpha\dot{s}s^{-3}\textbf{q}\nabla_{\textbf{r}_{i}}U(\textbf{r}_{0}).\end{split}$
In this case, $\alpha$ has dimensions of inverse of time. This relation between $\nabla_{\dot{\textbf{r}}_{i}}\mathcal{D}$ and $\nabla_{\textbf{r}_{i}}U$ allows us to obtain the equation for the homographic parameters ($s(t),\theta(t)$) in the presence of dissipation.
Substituting the homographic form $\textbf{r}_{h}(t)=\textbf{c}(t)+s(t)\textbf{Q}(\theta(t))(\textbf{r}_{0}-\textbf{c}_{0})$ (Eq. (17)) into the stacked equations of motion (12), and using that $\ddot{\textbf{c}}=0$, with identities (21) and multiplying by $\textbf{Q}^{-1}$, gives
@@
$\nabla_{\textbf{r}_{0}}U(\textbf{r}_{0})+\lambda\textbf{M}\left(\textbf{r}_{0}-\textbf{c}_{0}\right)=0,$
noticing that $\lambda=\frac{U(\textbf{r})}{I(\textbf{r})}<0$. Simultaneously, since J is a right-angle rotation, it renders $\textbf{J}(\textbf{r}_{0}-\textbf{c}_{0})$ orthogonal to $(\textbf{r}_{0}-\textbf{c}_{0})$. This splits equation (22) into terms parallel and orthogonal to $(\textbf{r}_{0}-\textbf{c}_{0})$, yielding a system of equations for the coefficients
$\begin{cases}\ddot{s}-s\dot{\theta}^{2}=\lambda\left(s^{-2}+\alpha\dot{s}s^{-3}\right)\\ s\ddot{\theta}+2\dot{s}\dot{\theta}=0\end{cases}$ (23)
Equation (23) is a dissipative Kepler-like system in the homographic parameters, with $s(t)$ playing the role of the radial coordinate, and $\theta(t)$ the angular position. Due to this equivalence, the next section treats the two-body problem under dissipation with $d=3$, providing solutions applicable both to the two-body case and to generic CCs for arbitrary $N$.
@@ 4 Two bodies
We describe the two-body system in polar coordinates $(r,\theta)$, where $v=\dot{r}$ denotes the radial velocity. We choose the mass unit such that the reduced mass is one, $m_{1}m_{2}/(m_{1}+m_{2})=1$. The equations of motion take the first-order form
@@
$\mathcal{H}=\frac{v^{2}}{2}+\frac{C^{2}}{2r^{2}}-\frac{\gamma}{r}\qquad\text{and}\qquad\mathcal{D}=\frac{\alpha}{2}\frac{\gamma}{r^{3}}v^{2},$ (25)
and the (conserved) normalized angular momentum CC is defined by
@@
It is convenient to define the non-dimensional parameter
$c=\frac{C}{\gamma^{2/3}\alpha^{1/3}}.$ (27)
For $\alpha>0$ and in convenient time and length scales, the dynamics depends only on $c$, see equation (42).
To understand the solutions of equation (24), we proceed to a Poincaré compactification of the phase space in a series of transformations, following Dumortier et al. [Dumortier_book] (Chapter 5.1). The idea relies on mapping the infinite phase space $(r,v)\in[0,\infty)\times\mathbb{R}$ into a sphere, and analyze the dynamics at infinity via stereographic charts (planar maps) covering the sphere. This allows us to understand the asymptotic behavior of solutions as they approach infinity in the original coordinates.
@@ 4.1 Poincaré compactification
$\begin{split}a^{c}_{\pm}=-\frac{\alpha\gamma}{2}\left(1\pm\sqrt{1-\frac{4C^{6}}{\gamma^{4}\alpha^{2}}}\right)=-\frac{\alpha\gamma}{2}\left(1\pm\sqrt{1-4c^{6}}\right),\end{split}$ (31)
@@
Hence, the equilibrium $(r_{e},v_{e})$ has an eigenvalue $a^{e}=-\gamma\alpha$ and a stable manifold coinciding with the axis $r=0$. The other eigenvalue is zero and the associated center-manifold $W_{e}^{c}$ is tangent to the vector $(\alpha,-1)$ (Fig. 1 (C)). Locally,
$W^{c}_{e}\approx\left\{v=\frac{C^{2}}{\gamma\alpha}-\frac{r}{\alpha}\right\},\quad\text{for }r>0\text{ small}.$
All orbits with initial conditions $v(0)<0$ and arbitrarily small $r(0)>0$ must approach $W^{c}_{e}$ before entering the circularization spiral, as shown in panel (C).
Figure 1: Phase space portraits for the regularized Kepler system (eq. 28) for $C=\gamma=1$. (A) The conservative case ($\alpha=0$) showing elliptic (blue), parabolic (green), and hyperbolic (orange) orbits. (B) The system with dissipation ($\alpha=0.5$) showing a captured orbit (blue), the weak-unstable-center manifold (red) from the equilibrium $(r_{e},v_{e})=(0,2)$, and an escaping orbit (orange); the green dot marks the circular equilibrium $(r_{c},v_{c})=(1,0)$. (C) Zoom of (B) near $(r_{e},v_{e})$; stable branches along $r=0$ are shown in blue.
@@
$(r,v)\longmapsto(y,x)\quad\text{where}\quad r=\frac{1}{y},\;\;v=\frac{x}{y},$
@@
In the new coordinates, equation (24) becomes
@@
$r_{\infty}\coloneqq(y_{\infty},x_{\infty})=(0,0).$
As the linear approximation of (32) at $(0,0)$ is null, $(y_{\infty},x_{\infty})$ is a degenerate equilibrium. One may note that all unbounded solutions of the Kepler problem ($\alpha=0$) are homoclinic to $(y_{\infty},x_{\infty})=(0,0)$ (see Fig. 2 (A)).
@@
To analyze the infinity we blow-up the vector-field at $(y,v)=(0,0)$ using [Dumortier_book, Sec. 3.3]
$(y,x)\longmapsto(p,q)\quad\text{where}\quad y=q^{2},\;\;x=pq^{3}.$
In the new coordinates the vector field (32), after a multiplication by $2q^{-3}$, becomes
$\begin{cases}\dot{q}=-qp\\ \dot{p}=-2\gamma+2C^{2}q^{2}-2\alpha\gamma pq^{3}+p^{2}\end{cases}$ (33)
The $p$-axis, which corresponds to $q^{2}=y=\frac{1}{r}=0$, is invariant. There are two hyperbolic singularities on this axis:
$(q_{\infty+},p_{\infty+})=\left(0,\sqrt{2\gamma}\right),\quad\text{and}\quad(q_{\infty-},p_{\infty-})=\left(0,-\sqrt{2\gamma}\right),$
connected by a solution from $(q_{\infty+},p_{\infty+})$ to $(q_{\infty-},p_{\infty-})$. The whole axis $p$ comes from the blow-up of the equilibrium $(y_{\infty},x_{\infty})$.
The equilibrium $(q_{\infty+},p_{\infty+})$ has a stable manifold $\Gamma_{s}$ that, up to linear approximation, is perpendicular to the $p$-axis: $\Gamma_{s}\approx\{q,p=\sqrt{2\gamma}\}$ for $|q|$ small. Similarly, $(q_{\infty-},p_{\infty-})$ has an unstable manifold $\Gamma_{u}$ that, up to linear approximation, is perpendicular to the $p$-axis: $\Gamma_{u}\approx\{q,p=-\sqrt{2\gamma}\}$ for $|q|$ small.
@@
$\Gamma_{s}=\sqrt{2\gamma}\,y^{\frac{3}{2}}\,,\qquad\Gamma_{u}=-\sqrt{2\gamma}\,y^{\frac{3}{2}}\,,\qquad y>0,$ (34)
@@
For $\alpha=0$, the vector field (32) has the first integral $\mathcal{H}=\frac{y^{3}\left(C^{2}y-2\gamma\right)+x^{2}}{2y^{2}}$ (from eq. 25), and both characteristic curves $\Gamma_{u}$ and $\Gamma_{s}$ are determined by the equation $\mathcal{H}=0$. One can check that this leads to the approximation in equation (34).
In Fig. 2 we show the local structure of the orbits of the vector field (32) for some values of $\alpha$. The stable curve $\Gamma_{s}$ approaching $(y_{\infty},x_{\infty})$ for positive time, and the curve $\Gamma_{u}$ approaching it for negative time divide the phase space xxyy into three sectors: a hyperbolic sector $S_{h}$, an attracting sector $S_{a}$, and a repelling sector $S_{r}$ (see Fig. 2 (C)).
In the Kepler problem, orbits with positive energy are asymptotically free, with $\mathcal{H}\to\frac{v^{2}}{2}$ as $t\to\pm\infty$. We define scattering orbits, in the dissipative or conservative scenarios, as those for which the following two limits exist
@@
In order to compute the asymptotic shape of a scattering orbit close to a point at infinity, one can use that
$\lim_{t\to\pm\infty}\mathcal{D}=\lim_{t\to\pm\infty}\frac{\alpha}{2}\frac{\gamma}{r^{3}}v^{2}=0$
to neglect dissipation of energy and solve
$\mathcal{H}=\frac{y^{3}\left(C^{2}y-2\gamma\right)+x^{2}}{2y^{2}}=\frac{v^{2}}{2},\quad\text{or}\quad y^{3}\left(C^{2}y-2\gamma\right)+x^{2}=v^{2}y^{2},$
@@
$x=\pm vy+\mathcal{O}(y^{3/2}).$
Therefore, a scattering orbit leaves the point at infinity tangent to the line $x=-v_{\psi}y$ through the repelling sector $S_{r}$, and arrives at the same point tangent to the line $x=v_{\omega}y$ through the attracting sector $S_{a}$ (see Fig. 2 (C)). In the presence of dissipation, $|v_{\omega}|<|v_{\psi}|$.
@@
$(r,v)\mapsto(w,z)\quad\text{where}\quad w=\frac{r}{v},\quad z=\frac{1}{v}$ (35)
In these coordinates, $z=0$ corresponds to $v=\pm\infty$. At first we will be interested on the limit as $v\to\infty$ that corresponds to the quadrant $w>0$, $z>0$.
After multiplying by $w^{3}$, the vector field (24) in the new chart becomes
@@
$v_{\infty}:=(w_{v\infty},z_{v\infty})=(0,0).$ (37)
The circular equilibrium $(r_{c},v_{c})$ is mapped to infinity in this chart.
In Fig. 3 we show phase portraits for the system with and without dissipation. The correctness of this numerically computed phase portrait is corroborated by the following analysis.
We blow up the vector field at $(w_{v\infty},z_{v\infty})=(0,0)$ using [Dumortier_book, Sec. 3.3]
$(w,z)\longmapsto(p,q)\quad\text{where}\quad w=pq^{3},\;\;z=q^{2},$ (38)
to obtain, after division by $q^{6}/2$,
[…]
In the region $p\geq 0$, $q\geq 0$, this equation has three equilibria: $(p,q)=(0,0)$, $(p_{e},q_{e})=(0,\sqrt{\frac{\alpha\gamma}{C^{2}}})$, which corresponds to the ejection point, and a new one $(p,q)=(\sqrt{\frac{\alpha\gamma}{2}},0)$.
Both equilibria $(p,q)=(0,0)$ and $(p,q)=\left(\sqrt{\frac{\alpha\gamma}{2}},0\right)$ are hyperbolic. The equilibrium $(p,q)=(0,0)$ is a saddle, with its unstable manifold lying on the $q$-axis and its stable manifold along the $p$-axis. The equilibrium $(p,q)=\left(\sqrt{\frac{\alpha\gamma}{2}},0\right)$ is a source with strong unstable manifold on the $p$-axis (eigenvalue $2\alpha\gamma$) and weak unstable manifold locally given by $W^{ws}_{\infty}\approx\{p=\sqrt{\frac{\alpha\gamma}{2}},q\}$ for $|q|$ small (eigenvalue $\alpha\gamma$).
In the variables $(w,z)$, the positive branch of $W^{ws}_{\infty}$ is given by
$W^{ws}_{\infty}\approx\left\{w=\sqrt{\frac{\alpha\gamma}{2}}z^{3/2}\right\}\,,\qquad z>0\ \ \text{small},$
while in the original $(r,v)$ variables it corresponds to
$W^{ws}_{\infty}\approx\left\{v=\frac{\alpha\gamma}{2}\frac{1}{r^{2}}\right\}\,,\qquad r>0\ \ \text{small}.$